OpenStudy (anonymous):

How many different pairings (2 teams playing each other) are possible for the first round of the 64 teams in the NCAA Basketball Tournament? (without seeding)

OpenStudy (kropot72):

\[64C2=\frac{64!}{2!(64-2)!}=\]

OpenStudy (kinggeorge):

Forgive me. Kropot is correct. We should be using combinations and not permutations.

OpenStudy (kropot72):

Answer is:\[\frac{64\times 63}{2}\]